OpenStudy (anonymous):

A bank credit card charges interest at the rate of 24% per year, compounded monthly. If a senior in college charges $2,600 to pay for college expenses, and intends to pay it in one year, what will she have to pay? (Round your answer to the nearest cent.)

OpenStudy (anonymous):

1year * 24%/year = 24% $2600(1+24%) = $2600*(1.24) = $3224

OpenStudy (anonymous):

thats what we tried but it says wrong answer...

OpenStudy (anonymous):

I think you have to use the formula a=Pe^rt

OpenStudy (anonymous):

But that is compounded continuously?

OpenStudy (anonymous):

sorry formula was A=P(1+r/k)^kt got the correct answer... thanks for your help!! correct answer was 3297.43